4. Valle de Vallibierna

’IBÓN DE LLOSÁS’ (LITTLE LAKE OF LLOSÁS) 1 hour 30 min.

The departure point is the refuge of Coronat at the end of the trail of Vallibierna, which has to be taken near the camping area close to the Paso Nuevo dam.

Follow the remains of a trail that goes to the valley on the right. The trail joins a well-planned track, which goes to the Plleta de Llosás, where you will get in about 30 minutes. Then follow the track on the left of the slope till you get to the Ibón de Llosás (2,490 m. high), which is just under the peaks Aneto, Tempestades and Rusell.

Follow the right bank of the ibón, come up the first ravine and you will get to the Ibonet de Llosás in 10 minutes.